• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Nerses

    Log entry time 23:07:47 on April 11, 2006

    Entry number 170488

    This entry is a followup to: 170486

    keyword=Back to production

    with run 3888 we back to production

    Beam positions, prescale values, event rate, DC HV are checked and are correct. Feedback is on, Happex and Compton are runing, raster and qasy are looking fine.

    Checked with onlineGUI: marker detector data were changed. I mean both TDC and ADC.




    Figure 1 Beam position and current



    Figure 2 BH Marker detectors



    Figure 3 Qasy



    Figure 4 raster